x^2p^²-2xyp+2y^2-x^2=0
Using Claratiurs method
1
Expert's answer
2020-08-26T15:59:38-0400

The equation above is not reducible to the Clairaut's Form as it can't be put in the form

"y = xp +f(p)"


To solve this equation, simple use "x^2p^2-2xyp+y^2-x^2 = 0"

Factorizing the above equation.

we get, "y = xp - x" and "y = xp +x"


Solving the equations,

"y = x(\frac{dy}{dx} - 1) \implies \frac{dy}{dx} - \frac{y}{x} = 1"


I.F. "e^{- \int \frac{dx}{x} } = \frac{1}{x}"

Then,

"\frac{1}{x} y = \int \frac{1}{x} dx"


"\frac{1}{x} y = lnx + ln C"


"y = x(lnx + ln C) = xln(Cx)"


Similarly for "y = x(p+1) \implies \frac{dy}{dx} - \frac{y}{x} = -1"


I.F. "e^{- \int \frac{dx}{x} } = \frac{1}{x}"

"\frac{1}{x} y = \int -\frac{1}{x} dx"


"y = -x(lnx + ln C) = -xln(Cx)"